OpenStudy (anonymous):

How would i find the solution to this equation? 64^(2 – x) = 4^4x

OpenStudy (anonymous):

2log64 - xlog64 = 4xlog4

OpenStudy (zepp):

You can try to make both sides of the equation of the same base

OpenStudy (zepp):

Using logarithms to solve this is unnecessary, you can apply this awesome rule here: \[a^b=a^c\\\text{Since both exponential expressions are in the same base, a, we can simply say that} \\b=c\]

OpenStudy (anonymous):

so 2-x=4x?

OpenStudy (anonymous):

yea totally

OpenStudy (zepp):

This rule can't be applied unless the bases are the same!

OpenStudy (zepp):

Here, let's put them on the same base, say, 4, because 64 is 4^3 \[\large 64^{2 – x} = 4^{4x}\\\large (4^3)^{2-x}=4^{4x}\]Apply the rule \(\large (a^{b})^{c}=a^{bc}\)\[\LARGE 4^{3(2-x)}=4^{4x}\]

OpenStudy (zepp):

Since they are on the same base, solve for 3(2-x)=4x and you are done :)

OpenStudy (anonymous):

would it be x=6 ?

OpenStudy (zepp):

Nope, try again :D

OpenStudy (zepp):

Distribute the 3 into the parenthses 3*2-3*x = 4x 6-3x=4x

OpenStudy (anonymous):

x=.86

OpenStudy (anonymous):

?